%% BioMed_Central_Bib_Style_v1.01

\begin{thebibliography}{16}
% BibTex style file: bmc-mathphys.bst (version 2.1), 2014-07-24
\ifx \bisbn   \undefined \def \bisbn  #1{ISBN #1}\fi
\ifx \binits  \undefined \def \binits#1{#1}\fi
\ifx \bauthor  \undefined \def \bauthor#1{#1}\fi
\ifx \batitle  \undefined \def \batitle#1{#1}\fi
\ifx \bjtitle  \undefined \def \bjtitle#1{#1}\fi
\ifx \bvolume  \undefined \def \bvolume#1{\textbf{#1}}\fi
\ifx \byear  \undefined \def \byear#1{#1}\fi
\ifx \bissue  \undefined \def \bissue#1{#1}\fi
\ifx \bfpage  \undefined \def \bfpage#1{#1}\fi
\ifx \blpage  \undefined \def \blpage #1{#1}\fi
\ifx \burl  \undefined \def \burl#1{\textsf{#1}}\fi
\ifx \doiurl  \undefined \def \doiurl#1{\url{https://doi.org/#1}}\fi
\ifx \betal  \undefined \def \betal{\textit{et al.}}\fi
\ifx \binstitute  \undefined \def \binstitute#1{#1}\fi
\ifx \binstitutionaled  \undefined \def \binstitutionaled#1{#1}\fi
\ifx \bctitle  \undefined \def \bctitle#1{#1}\fi
\ifx \beditor  \undefined \def \beditor#1{#1}\fi
\ifx \bpublisher  \undefined \def \bpublisher#1{#1}\fi
\ifx \bbtitle  \undefined \def \bbtitle#1{#1}\fi
\ifx \bedition  \undefined \def \bedition#1{#1}\fi
\ifx \bseriesno  \undefined \def \bseriesno#1{#1}\fi
\ifx \blocation  \undefined \def \blocation#1{#1}\fi
\ifx \bsertitle  \undefined \def \bsertitle#1{#1}\fi
\ifx \bsnm \undefined \def \bsnm#1{#1}\fi
\ifx \bsuffix \undefined \def \bsuffix#1{#1}\fi
\ifx \bparticle \undefined \def \bparticle#1{#1}\fi
\ifx \barticle \undefined \def \barticle#1{#1}\fi
\bibcommenthead
\ifx \bconfdate \undefined \def \bconfdate #1{#1}\fi
\ifx \botherref \undefined \def \botherref #1{#1}\fi
\ifx \url \undefined \def \url#1{\textsf{#1}}\fi
\ifx \bchapter \undefined \def \bchapter#1{#1}\fi
\ifx \bbook \undefined \def \bbook#1{#1}\fi
\ifx \bcomment \undefined \def \bcomment#1{#1}\fi
\ifx \oauthor \undefined \def \oauthor#1{#1}\fi
\ifx \citeauthoryear \undefined \def \citeauthoryear#1{#1}\fi
\ifx \endbibitem  \undefined \def \endbibitem {}\fi
\ifx \bconflocation  \undefined \def \bconflocation#1{#1}\fi
\ifx \arxivurl  \undefined \def \arxivurl#1{\textsf{#1}}\fi
\csname PreBibitemsHook\endcsname

%%% 1
\bibitem[\protect\citeauthoryear{Press et~al.}{2021}]{press2021train}
\begin{bchapter}
\bauthor{\bsnm{Press}, \binits{O.}},
\bauthor{\bsnm{Smith}, \binits{N.}},
\bauthor{\bsnm{Lewis}, \binits{M.}}:
\bctitle{Train short, test long: Attention with linear biases enables input length extrapolation}.
In: \bbtitle{International Conference on Learning Representations}
(\byear{2021})
\end{bchapter}
\endbibitem

%%% 2
\bibitem[\protect\citeauthoryear{Vaswani et~al.}{2017}]{vaswani2017attention}
\begin{botherref}
\oauthor{\bsnm{Vaswani}, \binits{A.}},
\oauthor{\bsnm{Shazeer}, \binits{N.}},
\oauthor{\bsnm{Parmar}, \binits{N.}},
\oauthor{\bsnm{Uszkoreit}, \binits{J.}},
\oauthor{\bsnm{Jones}, \binits{L.}},
\oauthor{\bsnm{Gomez}, \binits{A.N.}},
\oauthor{\bsnm{Kaiser}, \binits{{\L}.}},
\oauthor{\bsnm{Polosukhin}, \binits{I.}}:
Attention is all you need.
Advances in neural information processing systems
\textbf{30}
(2017)
\end{botherref}
\endbibitem

%%% 3
\bibitem[\protect\citeauthoryear{Devlin et~al.}{2018}]{devlin2018bert}
\begin{botherref}
\oauthor{\bsnm{Devlin}, \binits{J.}},
\oauthor{\bsnm{Chang}, \binits{M.-W.}},
\oauthor{\bsnm{Lee}, \binits{K.}},
\oauthor{\bsnm{Toutanova}, \binits{K.}}:
Bert: Pre-training of deep bidirectional transformers for language understanding.
arXiv preprint arXiv:1810.04805
(2018)
\end{botherref}
\endbibitem

%%% 4
\bibitem[\protect\citeauthoryear{Brown et~al.}{2020}]{brown2020language}
\begin{barticle}
\bauthor{\bsnm{Brown}, \binits{T.}},
\bauthor{\bsnm{Mann}, \binits{B.}},
\bauthor{\bsnm{Ryder}, \binits{N.}},
\bauthor{\bsnm{Subbiah}, \binits{M.}},
\bauthor{\bsnm{Kaplan}, \binits{J.D.}},
\bauthor{\bsnm{Dhariwal}, \binits{P.}},
\bauthor{\bsnm{Neelakantan}, \binits{A.}},
\bauthor{\bsnm{Shyam}, \binits{P.}},
\bauthor{\bsnm{Sastry}, \binits{G.}},
\bauthor{\bsnm{Askell}, \binits{A.}}, \betal:
\batitle{Language models are few-shot learners}.
\bjtitle{Advances in neural information processing systems}
\bvolume{33},
\bfpage{1877}--\blpage{1901}
(\byear{2020})
\end{barticle}
\endbibitem

%%% 5
\bibitem[\protect\citeauthoryear{Almazrouei et~al.}{2023}]{almazrouei2023falcon}
\begin{botherref}
\oauthor{\bsnm{Almazrouei}, \binits{E.}},
\oauthor{\bsnm{Alobeidli}, \binits{H.}},
\oauthor{\bsnm{Alshamsi}, \binits{A.}},
\oauthor{\bsnm{Cappelli}, \binits{A.}},
\oauthor{\bsnm{Cojocaru}, \binits{R.}},
\oauthor{\bsnm{Debbah}, \binits{M.}},
\oauthor{\bsnm{Goffinet}, \binits{E.}},
\oauthor{\bsnm{Heslow}, \binits{D.}},
\oauthor{\bsnm{Launay}, \binits{J.}},
\oauthor{\bsnm{Malartic}, \binits{Q.}}, et al.:
Falcon-40b: an open large language model with state-of-the-art performance.
Technical report,
Technical report, Technology Innovation Institute
(2023)
\end{botherref}
\endbibitem

%%% 6
\bibitem[\protect\citeauthoryear{Touvron et~al.}{2023}]{touvron2023llama}
\begin{botherref}
\oauthor{\bsnm{Touvron}, \binits{H.}},
\oauthor{\bsnm{Lavril}, \binits{T.}},
\oauthor{\bsnm{Izacard}, \binits{G.}},
\oauthor{\bsnm{Martinet}, \binits{X.}},
\oauthor{\bsnm{Lachaux}, \binits{M.-A.}},
\oauthor{\bsnm{Lacroix}, \binits{T.}},
\oauthor{\bsnm{Rozi{\`e}re}, \binits{B.}},
\oauthor{\bsnm{Goyal}, \binits{N.}},
\oauthor{\bsnm{Hambro}, \binits{E.}},
\oauthor{\bsnm{Azhar}, \binits{F.}}, et al.:
Llama: Open and efficient foundation language models.
arXiv preprint arXiv:2302.13971
(2023)
\end{botherref}
\endbibitem

%%% 7
\bibitem[\protect\citeauthoryear{Dosovitskiy et~al.}{2021}]{dosovitskiy2021an}
\begin{bchapter}
\bauthor{\bsnm{Dosovitskiy}, \binits{A.}},
\bauthor{\bsnm{Beyer}, \binits{L.}},
\bauthor{\bsnm{Kolesnikov}, \binits{A.}},
\bauthor{\bsnm{Weissenborn}, \binits{D.}},
\bauthor{\bsnm{Zhai}, \binits{X.}},
\bauthor{\bsnm{Unterthiner}, \binits{T.}},
\bauthor{\bsnm{Dehghani}, \binits{M.}},
\bauthor{\bsnm{Minderer}, \binits{M.}},
\bauthor{\bsnm{Heigold}, \binits{G.}},
\bauthor{\bsnm{Gelly}, \binits{S.}},
\bauthor{\bsnm{Uszkoreit}, \binits{J.}},
\bauthor{\bsnm{Houlsby}, \binits{N.}}:
\bctitle{An image is worth 16x16 words: Transformers for image recognition at scale}.
In: \bbtitle{International Conference on Learning Representations}
(\byear{2021})
\end{bchapter}
\endbibitem

%%% 8
\bibitem[\protect\citeauthoryear{Oquab et~al.}{2023}]{oquab2023dinov2}
\begin{botherref}
\oauthor{\bsnm{Oquab}, \binits{M.}},
\oauthor{\bsnm{Darcet}, \binits{T.}},
\oauthor{\bsnm{Moutakanni}, \binits{T.}},
\oauthor{\bsnm{Vo}, \binits{H.}},
\oauthor{\bsnm{Szafraniec}, \binits{M.}},
\oauthor{\bsnm{Khalidov}, \binits{V.}},
\oauthor{\bsnm{Fernandez}, \binits{P.}},
\oauthor{\bsnm{Haziza}, \binits{D.}},
\oauthor{\bsnm{Massa}, \binits{F.}},
\oauthor{\bsnm{El-Nouby}, \binits{A.}}, et al.:
Dinov2: Learning robust visual features without supervision.
arXiv preprint arXiv:2304.07193
(2023)
\end{botherref}
\endbibitem

%%% 9
\bibitem[\protect\citeauthoryear{Chen et~al.}{2021}]{chen2021decision}
\begin{barticle}
\bauthor{\bsnm{Chen}, \binits{L.}},
\bauthor{\bsnm{Lu}, \binits{K.}},
\bauthor{\bsnm{Rajeswaran}, \binits{A.}},
\bauthor{\bsnm{Lee}, \binits{K.}},
\bauthor{\bsnm{Grover}, \binits{A.}},
\bauthor{\bsnm{Laskin}, \binits{M.}},
\bauthor{\bsnm{Abbeel}, \binits{P.}},
\bauthor{\bsnm{Srinivas}, \binits{A.}},
\bauthor{\bsnm{Mordatch}, \binits{I.}}:
\batitle{Decision transformer: Reinforcement learning via sequence modeling}.
\bjtitle{Advances in neural information processing systems}
\bvolume{34},
\bfpage{15084}--\blpage{15097}
(\byear{2021})
\end{barticle}
\endbibitem

%%% 10
\bibitem[\protect\citeauthoryear{Janner et~al.}{2021}]{janner2021offline}
\begin{barticle}
\bauthor{\bsnm{Janner}, \binits{M.}},
\bauthor{\bsnm{Li}, \binits{Q.}},
\bauthor{\bsnm{Levine}, \binits{S.}}:
\batitle{Offline reinforcement learning as one big sequence modeling problem}.
\bjtitle{Advances in neural information processing systems}
\bvolume{34},
\bfpage{1273}--\blpage{1286}
(\byear{2021})
\end{barticle}
\endbibitem

%%% 11
\bibitem[\protect\citeauthoryear{Shaw et~al.}{2018}]{shaw2018self}
\begin{bchapter}
\bauthor{\bsnm{Shaw}, \binits{P.}},
\bauthor{\bsnm{Uszkoreit}, \binits{J.}},
\bauthor{\bsnm{Vaswani}, \binits{A.}}:
\bctitle{Self-attention with relative position representations}.
In: \bbtitle{Proceedings of the 2018 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 2 (Short Papers)},
pp. \bfpage{464}--\blpage{468}
(\byear{2018})
\end{bchapter}
\endbibitem

%%% 12
\bibitem[\protect\citeauthoryear{Huang et~al.}{2020}]{huang2020improve}
\begin{botherref}
\oauthor{\bsnm{Huang}, \binits{Z.}},
\oauthor{\bsnm{Liang}, \binits{D.}},
\oauthor{\bsnm{Xu}, \binits{P.}},
\oauthor{\bsnm{Xiang}, \binits{B.}}:
Improve transformer models with better relative position embeddings.
Findings of the Association for Computational Linguistics: EMNLP 2020
(2020)
\end{botherref}
\endbibitem

%%% 13
\bibitem[\protect\citeauthoryear{Raffel et~al.}{2020}]{raffel2020exploring}
\begin{barticle}
\bauthor{\bsnm{Raffel}, \binits{C.}},
\bauthor{\bsnm{Shazeer}, \binits{N.}},
\bauthor{\bsnm{Roberts}, \binits{A.}},
\bauthor{\bsnm{Lee}, \binits{K.}},
\bauthor{\bsnm{Narang}, \binits{S.}},
\bauthor{\bsnm{Matena}, \binits{M.}},
\bauthor{\bsnm{Zhou}, \binits{Y.}},
\bauthor{\bsnm{Li}, \binits{W.}},
\bauthor{\bsnm{Liu}, \binits{P.J.}}:
\batitle{Exploring the limits of transfer learning with a unified text-to-text transformer}.
\bjtitle{The Journal of Machine Learning Research}
\bvolume{21}(\bissue{1}),
\bfpage{5485}--\blpage{5551}
(\byear{2020})
\end{barticle}
\endbibitem

%%% 14
\bibitem[\protect\citeauthoryear{Dao et~al.}{2022}]{dao2022flashattention}
\begin{barticle}
\bauthor{\bsnm{Dao}, \binits{T.}},
\bauthor{\bsnm{Fu}, \binits{D.}},
\bauthor{\bsnm{Ermon}, \binits{S.}},
\bauthor{\bsnm{Rudra}, \binits{A.}},
\bauthor{\bsnm{R{\'e}}, \binits{C.}}:
\batitle{Flashattention: Fast and memory-efficient exact attention with io-awareness}.
\bjtitle{Advances in Neural Information Processing Systems}
\bvolume{35},
\bfpage{16344}--\blpage{16359}
(\byear{2022})
\end{barticle}
\endbibitem

%%% 15
\bibitem[\protect\citeauthoryear{Dao}{2023}]{dao2023flashattention}
\begin{botherref}
\oauthor{\bsnm{Dao}, \binits{T.}}:
Flashattention-2: Faster attention with better parallelism and work partitioning.
arXiv preprint arXiv:2307.08691
(2023)
\end{botherref}
\endbibitem

%%% 16
\bibitem[\protect\citeauthoryear{Su et~al.}{2021}]{su2021roformer}
\begin{botherref}
\oauthor{\bsnm{Su}, \binits{J.}},
\oauthor{\bsnm{Lu}, \binits{Y.}},
\oauthor{\bsnm{Pan}, \binits{S.}},
\oauthor{\bsnm{Murtadha}, \binits{A.}},
\oauthor{\bsnm{Wen}, \binits{B.}},
\oauthor{\bsnm{Liu}, \binits{Y.}}:
Roformer: Enhanced transformer with rotary position embedding.
arXiv preprint arXiv:2104.09864
(2021)
\end{botherref}
\endbibitem

\end{thebibliography}

\documentclass[sn-mathphys,Numbered]{sn-jnl}

\usepackage{graphicx}%
\usepackage{multirow}%
\usepackage{amsmath,amssymb,amsfonts}%
\usepackage{mathrsfs}%
\usepackage[title]{appendix}%
\usepackage{xcolor}%
\usepackage{textcomp}%
\usepackage{manyfoot}%
\usepackage{booktabs}%
\usepackage{algorithm}%
\usepackage{algorithmicx}%
\usepackage{algpseudocode}%
\usepackage{listings}%

\begin{document}

\title[HyPE: Attention with Hyperbolic Biases for Relative Positional Encoding]{HyPE: Attention with Hyperbolic Biases for Relative Positional Encoding}


\author{\fnm{Giorgio} \sur{Angelotti}\footnote{Dr. Giorgio Angelotti, a physicist, data scientist, and machine learning engineer, earned his PhD in Computer Science from the University of Toulouse, France, in 2023.}}\email{giorgio.angelotti@isae-supaero.fr}

\affil{Independent Researcher}


\abstract{In Transformer-based architectures, the attention mechanism is inherently permutation-invariant with respect to the input sequence's tokens. To impose sequential order, token positions are typically encoded using a scheme with either fixed or learnable parameters. We introduce Hyperbolic Positional Encoding (HyPE), a novel method that utilizes hyperbolic functions' properties to encode tokens' relative positions. This approach biases the attention mechanism without the necessity of storing the $O(L^2)$ values of the mask, with $L$ being the length of the input sequence. HyPE leverages preliminary concatenation operations and matrix multiplications, facilitating the encoding of relative distances indirectly incorporating biases into the softmax computation. This design ensures compatibility with FlashAttention-2 and supports the gradient backpropagation for any potential learnable parameters within the encoding. We analytically demonstrate that, by careful hyperparameter selection, HyPE can approximate the attention bias of ALiBi \citep{press2021train}, thereby offering promising generalization capabilities for contexts extending beyond the lengths encountered during pretraining. The experimental evaluation of HyPE is proposed as a direction for future research.}

\keywords{relative positional encoding, attention bias, large context, transformer}

\maketitle

\section{Introduction}
Transformer-based networks \citep{vaswani2017attention} have become pervasive across various fields, including Natural Language Processing \citep{devlin2018bert, brown2020language, almazrouei2023falcon, touvron2023llama}, Computer Vision \citep{dosovitskiy2021an, oquab2023dinov2}, and Offline Reinforcement Learning \citep{chen2021decision,janner2021offline}. Within Transformer-based architectures, the attention mechanism in the Transformer block hones in on the interrelations between token pairs in the input sequence. However, this attention mechanism exhibits permutation invariance, which implies that Transformers do not inherently discern the absolute and relative positions of tokens in a sequence. To remedy this, Positional Encoding—employing either fixed or learnable parameters—is integrated into the architecture \citep{vaswani2017attention,shaw2018self}, thereby augmenting the effectiveness of Transformers. Nevertheless, these encoding strategies introduce limitations to their versatility, particularly in their ability to adapt. Specifically, learnable positional encodings, tailored to the length of the input sequence, become fixed upon training completion, in parallel with the network weights. This inflexibility poses significant challenges when dealing with sequences at testing time that surpass the lengths encountered during training \citep{press2021train}.

Subsequent sections will delve into the integration of positional encoding in Transformer-based architectures, highlighting the existing limitations and providing the rationale for the innovations presented in this study.

\section{Positional Encoding}
Let us first recap what is computed by the attention mechanism.

Let $X \in \mathbb{R}^{L\times d}$ represent the sequence of embedded tokens, with $Q, K, V \in \mathbb{R}^{L\times d}$ denoting the query, key, and value matrices, respectively, and $W^{Q}, W^{K}, W^{V} \in \mathbb{R}^{d \times d}$ being the corresponding linear operators. The attention mechanism's output (for a single head) is then calculated as:
\begin{equation}
\label{eq:attention}
O = \textrm{softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right)V.
\end{equation}
\citet{vaswani2017attention} introduced positional information into Transformer models by superimposing a sinusoidal signal—dependent on the token's absolute position—onto the $d$-dimensional embedding resulting from the initial embedding layer. This sinusoidal bias represents the token's position within the sequence through a frequency that varies with its position. This bias $a^{\textrm{Sin}} \in \mathbb{R}^{L \times d}$ is usually directly added to the embedding before computing the queries, keys and values: $X \leftarrow X + a^{\textrm{Sin}}$.

In pursuit of better generalization for larger contexts and accommodating sequences beyond pre-training lengths, it is argued that \textit{relative distances} should be prioritized over \textit{absolute positions} within the encoded representation. To this end, \citet{shaw2018self} proposed Relative Positional Representations. The work in \citep{shaw2018self} suggests modifies keys and values by summing to them learnable embeddings, whose paramaters are adjusted based on the relative positions of tokens. Following a similar approach, \citet{huang2020improve} expanded this method to absolute distance encoding and vector embeddings. These approaches, in spite of their performances, are computational expensive. In the T5 model \citep{raffel2020exploring}, an attention bias $a^{\textrm{T5}} \in \mathbb{R}^{L\times L}$—a function of the relative distances—is integrated into the attention computation to encode the relative positions.
Despite their adaptivity, such relative positional encodings necessitate the storage of the full $O(L^2)$ mask, which becomes impractical for extensive context windows. Moreover, recent algorithmic advances enabling $O(L)$ computation of self-attention do not fully accommodate the bias addition before the softmax \citep{dao2022flashattention, dao2023flashattention}. To date (30 October 2023), the backward propagation of gradients for attention bias parameters still faces stability issues in FlashAttention-1\footnote{Reference to the official Triton implementation: \url{https://github.com/Dao-AILab/flash-attention/blob/83aef842beec1037eb8c1d9c3ef3ed8aae80b091/flash_attn/flash_attn_triton.py}} \citep{dao2022flashattention}, while FlashAttention-2\footnote{Reference to the official implementation: \url{https://github.com/Dao-AILab/flash-attention/blob/83aef842beec1037eb8c1d9c3ef3ed8aae80b091/flash_attn/flash_attn_interface.py}}\citep{dao2023flashattention} does not yet support attention biases.

Rotary Positional Encoding (RoPE)\citep{su2021roformer} includes both absolute positions and relative distances. RoPE uses a block diagonal rotation matrix to multiply the $W^{\{Q,K\}}$ weights. This encodes the position by rotating vectors created from the stacked embeddings of consecutive queries and keys. The varying angle of rotation with sequence position ensures the encoding of absolute positions, while the relative position is captured through the vector construction from consecutive embeddings.
Since RoPE is injected through matmul operations and does not leverage learned parameters, it allows for faster training and inference than T5.

Nevertheless, all the presented method does not seem to generalize efficiently when the model is evaluated on extrapolation of sequences of length $L_{\textrm{Extra}} > L$ used for pre-training \citep{press2021train}. To tackle this issue, the Attention with Linear Biases (ALiBi) paradigm \citep{press2021train} introduces a static, non-learned linear relative positional bias directly into the softmax argument, efficiently scaling Transformer architectures for extrapolation on longer than training context windows. However, ALiBi still requires storing the $O(L^2)$ mask, which while compatible with FlashAttention-1, is not yet accommodated by FlashAttention-2.

Our research seeks to answer:

\noindent\emph{Can we devise a method that mirrors ALiBi's effectiveness, incorporates learnable parameters, and harnesses FlashAttention-2's full capabilities?}

We introduce Hyperbolic Positional Encoding (HyPE), an innovative approach that can be mathematically nearly equivalent to ALiBi. Instead of adding a relative distance bias to the softmax term, HyPE utilizes hyperbolic functions and matrix multiplication properties to incorporate the information into the $QV^{T}$ computation. This method integrates seamlessly into any Transformer architecture leveraging FlashAttention, eliminating the need to store the full $O(L^2)$ mask by \textit{indirectly} calculating it during matrix multiplication.

The forthcoming section provides a concise overview of hyperbolic functions and our principal contribution.
\vspace{-0.1cm}
\section{Hyperbolic Relative Positional Encoding}
Let us examine the hyperbolic sine function:
\begin{equation}
\label{eq:sinh}
2\sinh{(x)} = e^{x}-e^{-x},
\end{equation}
which can be expressed as truncated series
\begin{equation}
\label{eq:approx}
\sinh{(x)} \approx x + O(x^{3}) \quad \text{for} \quad |x| \ll 1.
\end{equation}
From Equation \ref{eq:sinh} follows that
\begin{equation}
\label{eq:sum}
2\sinh{(x-y)} = e^{x-y} - e^{y-x}.
\end{equation}
This trivial identity will be fundamental for the ensuing discussion.

Define $a ^ {\textrm{HyPE}} \in \mathbb{R}^{L \times L}$ such that
\begin{equation}
a ^ {\textrm{HyPE}}_{i,j} = -\tau \sinh{(\mu(j-i))}.
\end{equation}
Our goal is to efficiently integrate the hyperbolic positional bias $a^{\textrm{HyPE}}$ into the softmax function's argument in Equation \ref{eq:attention}.

Let $\eta^{\{Q,K\}} \in \mathbb{R}^{L\times 2}$ be:
\begin{equation}
\eta^{Q}_{i,j} = \dfrac{\tau\sqrt{d}}{2}e^{(-1)^{j+1}\mu i}
\end{equation}
\begin{equation}
\eta^{K}_{i,j} = (-1)^{j+1}e^{(-1)^{j}\mu i}
\end{equation}
with indices $i \in [0,\dots,L-1]$, $j \in [0,1]$ and parameters $(\mu,\tau) \in \mathbb{R}^2$. Here, $\mu$ represents the slope and $\tau$ the amplitude.

\noindent By concatenating $\eta^{\{Q,K\}}$ with $Q$ and $K$ respectively:
\begin{equation}
\label{eq:concat1}
\hat{Q} = \textrm{concat}(Q, \eta^{Q}),
\end{equation}
\begin{equation}
\label{eq:concat2}
\hat{K} = \textrm{concat}(K, \eta^{K}),
\end{equation}
we can express the product $\hat{Q}\hat{K}^{T}$ as:
\begin{equation}
\hat{Q}\hat{K}^{T} = QK^{T} + \sqrt{d} a ^ {\textrm{HyPE}}.
\end{equation}
To illustrate, consider the explicit matrix multiplication:
\begin{equation}
(\hat{Q}\hat{V}^{T})_{i,l} = \sum_{j=0}^{d+1}\hat{Q}_{i,j}\hat{V}_{l,j} = \sum_{j=0}^{d-1}\hat{Q}_{i,j}\hat{V}_{l,j} + \sum_{j=d}^{d+1}\hat{Q}_{i,j}\hat{V}_{l,j} = \sum_{j=0}^{d-1}Q_{i,j}V_{l,j}+ \sum_{j=0}^{1}\eta^{Q}_{i,j}\eta^{K}_{l,j}.
\end{equation}
By utilizing the earlier identity (Equation \ref{eq:sum}):
\begin{equation}
\sum_{j=1}^{2}\eta^{Q}_{i,j}\eta^{K}_{l,j} = \dfrac{\tau \sqrt{d}}{2}\left(-e^{\mu(l-i)}+e^{\mu(i-l)}\right) = -\tau \sqrt{d} \sinh{ (\mu(l-i))} = -\sqrt{d} a^{\textrm{HyPE}}_{i,l}.
\end{equation}
Subsequently, the softmax function can be redefined:
\begin{equation}
\hat{O} = \textrm{softmax} \left( \dfrac{\hat{Q}\hat{K}^{T}}{\sqrt{d}} \right)V = \textrm{softmax} \left( \dfrac{QK^{T}}{\sqrt{d}} + a^{\textrm{HyPE}}\right)V.
\end{equation}

\noindent In this fashion, the $a^{\textrm{HyPE}}$ bias is incorporated without explicit storage, calculated indirectly during matrix multiplication.

For effective tensor partitioning across devices, $\eta^{\{Q,K\}}$ can be stacked $n$ times, resulting in a dimension of $d+2n$ that is divisible along the parallelism dimension.

\subsection{Approximate Equivalence with ALiBi}
Using the truncated series expression (Equation \ref{eq:approx}), we deduce that for $\tau = 1$ and $\mu = m < \frac{1}{L}$, where $m$ is ALiBi's slope,
\begin{equation}
a^{\textrm{HyPE}}(\mu, \tau=1)_{i,j} \approx  a^{\textrm{ALiBi}}(m=\mu)_{i,j} + O\left((j-i)^{3}\mu^{3}\right) \quad \text{for} \quad \lvert (j-i) \mu \rvert \ll 1.
\end{equation}
This result suggests that for an accurate selection of (hyper)parameters HyPE can approximate ALiBi, while resorting to FlashAttention-2.
\subsection{Selection of (hyper)parameters}
Given the exponential divergence of $\sinh(x)$ for large $x$, we suggest setting $\mu < L^{-1}_{\textrm{Extra}}$, where $L_{\textrm{Extra}}$ exceeds the pre-training length $L$.

The amplitude $\tau$ may be static or dynamically adjusted through gradient descent.
\subsection{Multi-Head Attention Implementation}
HyPE accommodates multi-head attention by defining head-specific $\eta_{h}^{\{Q,K\}}$ and applying the concatenation steps. In this multi-head context, each head can have a unique slope $\mu_{h}$, and $\tau_{h}$ could be a learnable amplitude.

\subsection{Extending HyPE to Images and Multidimensional Data}
HyPE is versatile and can be extended to process multidimensional data, such as images. For instance, consider a 3D image structured as a parallelepiped with dimensions $(L_{x}, L_{y}, L_{z})$. Here, each voxel represents a token, identifiable by its 3D coordinates $(i,j,k)$.
These voxels are fed into the Transformer model as a flattened sequence of tokens with a total count of $L_{x}L_{y}L_{z}$.

To adapt HyPE for this context, one must carefully align the $\eta_{{x,y,z}}^{\{Q,K\}}$, matrices, ensuring that each token's dimensional indices are accurately represented.   It is crucial to verify that the $\eta$ matrices relative to the same dimension corresponding to the same dimension interact correctly during the matrix multiplication process. This precise arrangement allows the computation of attention as follows:
\begin{equation}
\hat{O} = \textrm{softmax} \left(\dfrac{QK^{T}}{\sqrt{d}} + a^{\textrm{HyPE}}_{x} + a^{\textrm{HyPE}}_{y} + a^{\textrm{HyPE}}_{z}\right)V.
\end{equation}
The same reasoning can be extended to input of any dimension.

\section{Experimental Evaluation}

The authors, as Independent Researchers with limited access to computational resources, have deferred the experimental evaluation of the algorithm to future work. We posit that a partial evaluation, constrained by computational limitations, might lead to an incomplete analysis that might not fully represent the merits of the proposed approach. Nonetheless, we invite researchers worldwide to apply the technique within their own computational environments and contribute to its empirical evaluation.
The primary challenges in implementing this approach may lie in the multiplication of low-precision values and the allocation of memory.
\section{Conclusion}

This study tackled the challenge of devising a Relative Positional Encoding that aligns with the results of ALiBi while leveraging the advancements in FlashAttention technology. Our development of Hyperbolic Positional Encoding (HyPE) requires storing only $O(4Lh)$ values—where $h$ represents the number of attention heads—and is capable of implicitly integrating a hyperbolic attention bias into the softmax argument of the attention mechanism via matrix multiplication of altered queries and keys. This process necessitates an additional memory allocation step for the queries before the attention computation, which can then be efficiently performed using FlashAttention. While the parameters of HyPE have the potential to be learnable, we demonstrate that with appropriate fixed values, it can approximate the attention bias characteristic of ALiBi. We also show that HyPE can incorporate relative positions for input of any dimension, e.g. images. Comprehensive experimental evaluation remains an objective for future work.
\backmatter

\bibliography{sn-bibliography}
\end{document}